The Outdoor Shop buy tents for $264 less 25% for buying more than 20 tents. The store operates on a markup of 33 1/3% of the sale price and advertises that all merchandise sold at a discount of 20% of the regular selling price. What is the regular selling price of the tents?
Do not understand how to break this down.

Answer by stanbon(75887)

The Outdoor Shop buy tents for $264 less 25% for buying more than 20 tents. The store operates on a markup of 33 1/3% of the sale price and advertises that all merchandise sold at a discount of 20% of the regular selling price. What is the regular selling price of the tents?
-------
Cost of the tents if over 20 purchased:: 0.75*264 = 198 per tent
--------
Markup price 1.33333*198 = 130.66
---------
Discounted price = 0.80*130.66 = $104.53
---------------
Cheers,
Stan H.
---------------